HUBBLE SPACE TELESCOPE DAILY REPORT #3089 PERIOD COVERED: 0000Z (UTC) 04/05/02 - 0000Z (UTC) 04/08/02 Daily Status Report as of 098/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.2 Completed Ten Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.3 Completed WF/PC-2 9039 (Getting to the Core of The Matter: The Inner Region of M15) The WF/PC-2 was used to measure the velocity dispersion profile within 15 arcseconds of the core of M15 to determine whether the mass distribution in this region is best fit by the presence of a Black Hole or by a more extended distribution of mass. Three models of the core region will be tested {1} a massive Black Hole, {2} a more extended distribution of neutron stars, or {3} is in the process of core collapse. All observations completed nominally. 1.4 Completed WF/PC-2 9060 (Photometry of a Statistically Significant Sample of Kuiper Belt Objects) The WF/PC-2 was used to propel the physical study of KBOs forward by performing accurate photometry at V, R, and I on a sample of up to 150 KBOs. The sample is made up of objects that will be observed at thermal infrared wavelengths by SIRTF and will be used with those data to derive the first accurate diameters and albedos for a large sample of KBOs. The observations completed nominally. 1.5 Completed Three Sets of STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.6 Completed Nine Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.7 Completed Four Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.8 Completed Thirty-two Sets of STIS/CCD 8910 (Sparse Field CTE Test-Internal {Cycle 10}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. This program needs special commanding to provide {a} off-center MSM positionings of some slits, and {b} the ability to read out with any amplifier {A, B, C, or D}. No problems were encountered. 1.9 Completed Five Sets of WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.10 Completed Two Sets of ACS/HRC/WFC 9028 (CCD Geometric Distortion) The Advanced Camera for Surveys (HRC and WFC) was used to measure the geometric distortion of the WFC and HRC detectors using multiple pointing observations of globular cluster 47 Tucanae , thus imaging moderately dense stellar fields. Centroid position of stars will be used to tie down the geometric distortion to about 0.2 pixels across the field of each detector. These high quality baseline measurements of the distortion will be ideal for diagnosing any temporal variations in later calibration campaigns. No problems were encountered. 1.11 Completed Three Sets of FGS/1/2 8995 (Verification of Operational FGS Calibrations) Fine Guidance Sensors #1 and #2 were used to monitor the characteristics of HST's Fine Guidance Sensors across the boundary of the servicing mission SM3B and compares the data with pre-SM3B baseline characterizations. For each FGS, stars from a selected field in the standard astrometric Galactic cluster M35 are observed in Position mode. This provides a measurement of their relative positions, which in turn provides an effective means to detect and monitor any changes in the FGS optical distortions and plate scales. These observations will obtain S-curves at the center of each FGS FOV. If changes larger than tolerance are observed, it will be necessary to activate the appropriate contingency proposal that will obtain data for new S-curves. These data will also be used to re-compute the optimal values of parameters, or K-factors, that are used in the acquisition and tracking of guide stars {and astrometry targets}. The proposal completed with no reported problems. 1.12 Completed ACS/HRC/WFC 9015 (CCD Image Quality Verification) The Advanced Camera for Surveys (HRC and WFC) was used to obtain a series of images to evaluate the point source image quality over the fields of view of the ACS HRC and WFC channels in normal imaging {non-coronagraphic} mode, after the corrector mechanisms have been used to optimize the image focus and symmetry. A moderate density open cluster, NGC-188, will be the target, since it is available year-round and has suitable density and magnitude distribution. Images will be obtained in the F502N and F625W filters, with 0.5 px dithers in each axis of the WFC to improve PSF sampling. Larger {~10 arcsec} POS-TARG offsets will be used to improve the sampling over the HRC field, while maintaining mean star separation {>5 arcsec} adequate for PSF evaluation to a reasonable radius. The proposal completed with no reported problems. 1.13 Completed Two Sets of STIS/CCD 9180 (Gamma-ray Burst Progenitors: Probing Their Environment) The Space Telescope Imaging Spectrograph (CCD) was used to perform a target of opportunity observation of gamma ray burster (GRB), GRB-020127. GRB astronomy is a field maturing at a phenomenal rate. Three important new observational and theoretical discoveries, formulated over the last twelve months, allow the proposer to address new, and in many cases, more sophisticated questions than could have been posed previously. These developments: the discovery of X-ray lines in GRB 991216; the observation that N_H as deduced from X-ray afterglow are one to two orders of magnitude larger than the dust extinction inferred from optical afterglow; and the growing realization that the afterglow emission may exhibit features of dust echoes, appear to offer unexpected and new diagnostics that will directly inform us about the progenitor, the circum-progenitor material and the immediate interstellar environs. There were no reported problems. 1.14 Completed Eight Sets of WF/PC-2 8952 (SM3B Flat Field Calibration) The WF/PC-2 was used to check for changes in flat field using external earth flat fields. No problems were encountered. 1.15 Completed STIS/CCD 9239 (Absolute Proper Motions of Nearby Dwarf Spheroidal Galaxies-Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to measure precise absolute proper motions for four dwarf spheroidal satellites of the Milky Way using spectroscopically-confirmed background QSOs to define a zero- velocity reference frame. The proposal completed with no reported problems. 1.16 Completed Two Sets of FGS/1 9034 (The Masses and Luminosities of Population II Stars) Fine Guidance Sensor #1 was used to observe the mass-luminosity relation {MLR} of Population II stars of which very little is currently known. With the advent of the Hipparcos Catalogue, improved distances to many spectroscopic binaries known to be Pop II systems are now available. After surveying the literature and making reasonable estimates of the secondary masses, we find 13 systems whose minimum separation should be larger than the resolution limit of FGS #1. The observations completed nominally. 1.17 Completed Two Sets of ACS/WFC/HRC 8992 (ACS Early Release Observations) The Advanced Camera for Surveys (WFC and HRC) was used to perform early release observations {ERO}, employing a variety of instruments to demonstrate the new capabilities. The observations (not all in this iteration) will include: 1} WFC observations of peculiar galaxies NGC 4676 {The Mice} and UGC 10214. This will be used to study young stellar clusters formed through the process of galaxy interaction. Any dark companion to UGC 10214 will also be searched for in sensitive ACS/WFC observations. 2} Emission line and continuum imaging of the Cone Nebula and M 17 with WFC to demonstrate the improved sensitivity and angular resolution of ACS. 3} Coronagraphic imaging of HD 141569, a main sequence star with circumstellar disk. 4} Coronagraphic imaging with the HRC of the nearby quasar 3C 273 to investigate the physical properties of the host galaxy and environments. There were no reported problems. 1.18 Completed STIS/CCD 9486 (What Excites LINERs: The Brilliant Case of NGC 3998) The Space Telescope Imaging Spectrograph (CCD) was used to observe LINER emission that is now known to occur in many galaxies at a variety of levels, with causes as diverse as old nuclear starbursts, cooling flows, accretion disks, and jets. The early prototypes, including NGC 3998, belong to a distinct class of AGN-like LINERs with the most luminous line emission and powerful nuclear radio emission. These LINERs have ionized gaseous disks on the scale of tens of parsecs, which have now been found to be in rotation about supermassive black holes. The disks are apparently the outer parts of energy- releasing accretion disks. Powerful new UV+optical emission line diagnostics have been developed to settle the long- standing debate over whether the gas is being photoionized or shock excited. There was no anomalous activity. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 24 Successful: 24 Scheduled Re-acquisitions: 22 Successful: 22 2.2 FHST Updates: Scheduled: 56 Successful: 56 2.3 Operations Notes: The NICMOS Cryo Cooler (NCS) cool down continues nominally. As of Sunday morning, the NICMOS Outlet Temperature (MNPNCOLT) was 91.99K (expected end point 76K), the NCC TA Inlet Temperature (MNPTAILT) was 75.18K (expected end point 68K), the CLI Temperature (MNPLINTT) was 79.50K (expected end point 71K), and the NICMOS Inlet Temperature (MNRNCILT) was 70.19K (expected endpoint 63K). During this weekend, the following operations notes were invoked: Limit Management during WF/PC-2 decontamination Change limits for MCE-2 reset detection Adjust STIS OTCCZ3BT limits An operations request was executed at 095/1814Z to inhibit MAMA-1 turn-on (ramp-up scheduled just after the period of this report). The TRSWCC relay command stack was re-ordered at 095/1924Z per an operations request. Using ROP IC-2, SSA transmitter #2 was turned on for the intervals 095/1409Z to 095/1438Z and from 096/0119Z to 096/0148Z. The STIS MCE-2 reset at 095/181430Z while outside any SAA interval and while the low voltage was on. MAMA-2 will be recovered at the next time that high voltage is commanded in the SMS. The SSR-1 EDAC error counter was cleared at 096/1253Z per ROP SR-1A. The mnemonic JWCK4P5 (WF/PC CEB Clock 4 +5v) flagged out-of-limits high at 096/150854Z with a value of 5.23447 volts. The parameter returned into normal ranges at 096/150954Z. HSTAR 8596 was written. The 486 engineering status buffer limits were adjusted for SESBSLD at 096/1615Z and at 096/1638Z, using ROP DF-18A. Per ROP SR-9A, the SSR-3 EDAC error counter was cleared at 097/1243Z. SI C&DH errors were reset at 097/1245Z and at 097/1710Z according to ROP NS-5. A TTR was written for a required re-transmit at 097/1702Z during a NSSC-1 load. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW